# Semantic Refinement with LLMs for Graph Representations

### v1

@@ 2.5 Memory-Guided Semantic Refinement
Given the history memory $\mathcal{B}^{(t)}$ at iteration $t$, DAS updates node semantics through a in-context refinement operator. This operator defines how the current description of each node is locally reshaped under task-aligned, in-graph references.
@@
For each node $v\in\mathcal{V}$, an exemplar set $\mathcal{S}_{v}^{(t)}\subset\mathcal{B}^{(t)}$ is first retrieved based on joint semantic–structural similarity and model stability. The large language model $\mathcal{M}$ is then applied as a conditional refinement operator $d_{v}^{(t+1)}=\mathcal{M}\big(d_{v}^{(t)},\mathcal{S}_{v}^{(t)}\big)$, where $d_{v}^{(t+1)}$ denotes the refined semantic description at iteration $t+1$.
The LLM is instructed to perform semantic reweighting and compression rather than knowledge expansion. Specifically, it reconstructs $d_{v}^{(t)}$ by emphasizing discriminative cues implicitly indicated by the exemplar set $\mathcal{S}_{v}^{(t)}$, while remaining faithful to the existing content. Since $\mathcal{S}_{v}^{(t)}$ is drawn from the same graph and filtered by the current classifier, the refinement is implicitly shaped by both structural context and task supervision.
@@
The refinement in Eq. (10) is applied to all nodes in parallel, yielding
$\mathcal{D}^{(t+1)}=\{d_{v}^{(t+1)}\mid v\in\mathcal{V}\}.$ (9)
These updated descriptions are re-encoded as node features for the next training stage of the fixed GNN, which in turn updates the predictive state used to construct the next memory $\mathcal{B}^{(t+1)}$.
Iterative Semantic Evolution.
Starting from the initialization $d_{v}^{(0)}=r_{v}$, the refinement proceeds as
$d_{v}^{(0)}\rightarrow d_{v}^{(1)}\rightarrow\cdots\rightarrow d_{v}^{(T)}.$ (10)
Early iterations primarily reflect coarse lexical and structural cues, while later iterations progressively concentrate on task-discriminative semantics shaped by model feedback. After $T$ iterations, the final descriptions $\{d_{v}^{(T)}\}$ are used for evaluation.
2.6 Why Iterative Refinement Works
To further explain why DAS benefits from iterative semantic refinement, we provide a theoretical analysis in Appendix B. Specifically, we formalize DAS as an alternating optimization process over model parameters and node semantics, where the model-conditioned memory induces a task-adaptive surrogate objective. We show that, under mild assumptions, each refinement iteration is guaranteed not to increase this global objective. This result explains why DAS can progressively improve node semantics without suffering from uncontrolled semantic drift, and clarifies that the refinement loop constitutes a principled optimization process rather than heuristic text rewriting.
@@ 3 Experiments
Nodes Edges Classes Graph Types
Cora 2,708 10,556 7 Text-attribute
Pubmed 19,717 88,648 3 Text-attribute
USA 1,190 28,388 4 Text-free
Europe 399 12,385 4 Text-free
Brazil 131 2,137 4 Text-free
Table 1: Dataset statistics used in our experiments.
@@ 3.1 Experimental Setup
We evaluate on five graphs following wang2024tans: two text-attributed citation networks, Cora and Pubmed, and three text-free airport networks, USA, Europe, and Brazil (Statistics are given in Table 1). For Cora/Pubmed, nodes are papers (title+abstract), edges are citations, and classes are research topics. For airports, nodes are airports, edges are flight connections, and classes correspond to activity levels (ribeiro2017struc2vec).
@@
Raw Text $76.97\pm 1.95$ $75.50\pm 2.03$ $70.78\pm 2.00$ $87.49\pm 0.54$ $87.20\pm 0.51$ $82.58\pm 0.38$ $4.17$
+ TAPE $76.50\pm 3.27$ $75.30\pm 1.92$ $71.06\pm 2.13$ $88.21\pm 0.62$ $87.80\pm 0.48$ $83.98\pm 0.59$ $3.67$
+ KEA $76.88\pm 1.73$ $75.74\pm 2.06$ $71.32\pm 2.51$ $88.10\pm 0.49$ $87.77\pm 0.50$ $85.33\pm 0.41$ $3.17$
+ TANS $76.27\pm 2.35$ $76.99\pm 2.02$ $73.64\pm 2.59$ $89.17\pm 1.17$ $7.98\pm 0.48 88$$8.84\pm 0.43$ $2.50$
+ DAS (Ours) $6.49\pm 2.57 77$$7.71\pm 3.19$ $74.30\pm 2.39$ $9.87\pm 0.55 88$$57\pm 0.64 88.$$8.95\pm 0.79$ $1.50$
@@
Random Walk $56.70\pm 2.47$ $56.11\pm 2.11$ $69.70\pm 14.34$ $55.71\pm 4.01$ $62.80\pm 3.01$ $68.40\pm 9.65$ $3.67$
TANS $55.56\pm 1.52$ $61.08\pm 2.71$ $65.73\pm 11.91$ $56.33\pm 5.73$ $65.81\pm 3.11$ $71.60\pm 10.66$ $2.67$
DAS (Ours) $\bm{56.80\pm 2.79}$ $60.01\pm 2.19$ $69.09\pm 10.65$ $\bm{59.38\pm 5.05}$ $\bm{67.23\pm 3.41}$ $\bm{75.19\pm 7.42}$ $\bm{1.50}$
@@
Baselines.
For text-attributed graphs, we compare Raw Feat. (bag-of-words/TF-IDF), Raw Text (use original text), TAPE (he2024harnessing), KEA(chen2024exploring), and TANS (wang2024tans). For text-free graphs, we compare hand-crafted topology features—Node Degree, Eigenvector (dwivedi2023benchmarking), Random Walk (dwivedi2022graph), and TANS wang2024tans. In all text-attributed baselines, generated texts are appended to the original node text and encoded by the same sentence encoder for fairness, while for text-free graphs, the generated texts are used directly as node descriptions.
Evaluation Protocol.
Unless otherwise specified, we focus on node classification with a GCN backbone (kipf2017semisupervised). We also report results with GAT (velickovic2018graph) and MLP in the text-attributed setting, following wang2024tans. For single-graph learning, we adopt the low-label / high-label splits: in the low-label regime, we use 20/30 nodes per class for train/valid on Cora/Pubmed (10/20 for Brazil); in the high-label regime, we use a 60/20/20 train/valid/test split. All reported numbers are averages over 30 random seeds with mean $\pm$ standard deviation, selecting models by the best validation accuracy. For the text encoder, we adopt MiniLM (wang2020minilm) unless otherwise noted.
3.2 Results on Text-Attributed Graphs
Table 2 reports node classification accuracy on text-attributed Cora and Pubmed under both low- and high-label settings, using GCN, GAT, and MLP backbones. On both of these datasets, augmenting Raw Text with existing LLM-based methods consistently improves performance over Raw Feat. and Raw Text across all backbones in both label regimes. Our method further improves upon the other methods in all configurations, consistently getting an accuracy higher than TANS for all backbones. These results indicate that structure-aware refinement is robust across architectures and scales of text-attributed citation graphs.
@@ 3.2 Results on Text-Attributed Graphs
3.3 Results on Text-Free Graphs
We evaluate DAS on three text-free airport graphs using GCN as the backbone. Table 3 reports results under both low- and high-label splits. In the low-label regime, structural baselines such as Node Degree, Eigenvector, and Random Walk outperform the one-hot feature baseline, confirming the importance of topology. DAS matches or surpasses these baselines on all three graphs, achieving the best accuracy on Europe and comparable performance on USA and Brazil. In the high-label regime, DAS yields clearer gains, achieving the strongest performance on USA and Brazil, and also outperforming all baselines on Europe. Notably, although no human-written text is available, DAS leverages structural cues to retrieve exemplars and progressively induces task-aligned node semantics through refinement. This structure-guided semantic evolution likely explains why gains on text-free graphs are larger than on text-attributed citation networks, where strong node texts already exist. Overall, these results indicate that iterative, structure-aware refinement is particularly effective when raw node features are absent and structural roles dominate.
3.4 Results under Distribution Shift
@@
4 Mechanism Analysis
This section analyzes the key mechanisms underlying DAS beyond aggregate performance. We focus on how iterative refinement and history-guided exemplar retrieval shape semantic evolution, and characterize when these mechanisms succeed or fail. We provide additional analysis in Appendix F
Figure 3: The impact of the number of iterations.
The Role of Iterative Refinement.
We analyze whether iterative refinement is necessary beyond a single-round semantic update. We vary the number of refinement iterations $T\in\{1,2,3\}$ on a text-attributed graph (Cora) and two text-free graphs (USA, Europe). Figure 3 reports the accuracy improvement over using raw node descriptions. Across all datasets, performance improves monotonically as $T$ increases, indicating that semantic refinement benefits from repeated feedback rather than one-shot rewriting. This trend suggests that successive rounds allow the model-conditioned memory to accumulate more reliable exemplars, which in turn guide the LLM toward increasingly task-aligned descriptions. The marginal gain from T=2T{=}2 to T=3T{=}3 is small, indicating that a limited number of iterations suffices to realize most of the benefit, while keeping LLM cost manageable.
@@ 4 Mechanism Analysis
Setting Cora USA
DAS (Ours) 89.31 $\pm$ 1.14 67.23 $\pm$ 3.41
@@
Random exemplars 89.06 $\pm$ 0.92 66.47 $\pm$ 3.24
Table 5: Ablation results on memory.
We examine the role of the model-conditioned memory in exemplar retrieval for semantic refinement. To isolate its effect, we compare DAS with three ablated variants: Random exemplar selection, Structure-only retrieval based solely on structural similarity, and Text-only retrieval based solely on semantic similarity. All variants use the same refinement procedure and prompt format.
@@
Results in Table 5 show that joint semantic–structural retrieval in DAS consistently outperforms all ablated variants on both Cora and USA. On Cora, text-only retrieval performs competitively, reflecting the strong semantic signal in raw node texts, while structure-only retrieval is weaker. In contrast, on the text-free USA graph, both text-only and structure-only retrieval degrade performance, indicating that neither modality alone is sufficient for stable refinement. Random exemplar selection performs worst in both cases. These results demonstrate that effective refinement requires task-relevant exemplars that are aligned in both semantic content and structural role. The joint retrieval mechanism is particularly critical when node semantics must be induced from topology, where relying on a single modality can introduce noisy or misleading in-context signals.
Why DAS Succeeds: Semantic Sharpening and Role Abstraction.
DAS is most effective when iterative refinement sharpens task-discriminative evidence already latent in the input representations. On text-attributed graphs, successful refinement makes class-consistent technical cues more explicit, leading to reduced predictive entropy and corrections toward the true label. On text-free graphs, DAS succeeds when raw topological statistics are reorganized into a coherent semantic role interpretation (e.g., distinguishing regionally embedded nodes from global connectors based on clustering and betweenness). In both cases, refinement aligns node semantics more closely with the structure–semantics regime of the graph, enabling the classifier to make more confident and accurate predictions. Representative examples are provided in Table 7 in the appendix.
When DAS Fails: Drift and Over-Confidence.
Failure cases reveal inherent limitations of LLM-based semantic refinement. One common failure mode is label drift, where rewrites improve fluency without introducing additional discriminative evidence, causing predictions to shift toward a semantically adjacent but incorrect class. Another failure mode is over-confidence, in which refinement reduces predictive entropy while the prediction itself remains incorrect. On text-free graphs, we additionally observe occasional attribute drift, where numeric structural attributes are subtly altered during rewriting, raising faithfulness concerns even when predictive accuracy improves. These failures highlight that effective refinement depends on maintaining a tight coupling between generated semantics and the underlying structural evidence. See Table 8 in the appendix for examples.
@@ 5 Related Work
Classical GNNs learn representations through local message passing, thereby encoding fixed inductive biases such as locality and homophily (kipf2017semisupervised; hamilton2017inductive; velickovic2018graph; xu2018how). Structure-only methods based on random walks or structural roles further impose predefined topological assumptions (perozzi2014deepwalk; grover2016node2vec; ribeiro2017struc2vec). While effective in specific regimes, these methods rely on fixed priors that do not adapt to heterogeneous structure–semantics distributions across graph domains.
Model-Centric Refinement.
@@
Most existing attempts to address structure–semantics heterogeneity remain model-centric. This includes strengthening GNN architectures with richer structural priors (maron2018invariant; murphy2019relational; jin2022raw), treating LLMs as direct graph reasoners via graph-to-text serialization (zhao2023graphtext; kong2024gofa; chen2024llaga), and introducing auxiliary models to enrich node representations with external semantic signals (yao2019textgcn; he2023harnessing; yang2021graphformers). Despite their diversity, these approaches all inject additional inductive biases from the model side, and the construction of node representations remains largely static once the model is specified.
Data-Centric Refinement and Our Perspective.
@@
Existing data-centric methods primarily focus on graph augmentation, structure learning, or pseudo-labeling for robustness and generalization (you2020graph; jin2020graph; chen2023label; he2023harnessing; chen2024exploring), rather than on adapting node semantics to task-specific structure–semantics regimes. In contrast, our work treats node semantics as task-adaptive variables and proposes a feedback-driven framework for iterative, structure-aware semantic refinement, providing a fundamentally different data-centric perspective for handling structure–semantics heterogeneity.
6 Conclusion

### v2

@@ 2.5 Memory-Guided Semantic Refinement
Given the memory $\mathcal{B}$, GES updates node semantics through an in-context refinement operator. This operator defines how the description of each node is locally reshaped under task-aligned, in-graph references.
@@
For each node $v\in\mathcal{V}$, an exemplar set $\mathcal{S}_{v}\subset\mathcal{B}$ is first retrieved based on joint semantic–structural similarity and model stability. The large language model $\mathcal{M}$ is then applied as a conditional refinement operator $d_{v}^{\prime}=\mathcal{M}(d_{v},\mathcal{S}_{v})$, where $d_{v}^{\prime}$ denotes the refined semantic description.
The LLM is instructed to perform semantic reweighting and compression rather than knowledge expansion. Specifically, it reconstructs $d_{v}$ by emphasizing discriminative cues implicitly indicated by the exemplar set $\mathcal{S}_{v}$, while remaining faithful to the existing content. Since $\mathcal{S}_{v}$ is drawn from the same graph and filtered by the current classifier, the refinement is implicitly shaped by both structural context and task supervision.
@@
The refinement is applied to all nodes in parallel, yielding $\mathcal{D}^{\prime}=\{d_{v}^{\prime}\mid v\in\mathcal{V}\}$. These refined descriptions are re-encoded as node features and a new GNN is trained on these encodings for evaluation.
@@ 3.1 Experimental Setup
We evaluate on five graphs following Wang et al. (2025a): two text-attributed citation networks, Cora and Pubmed, and three text-free airport networks, USA, Europe, and Brazil (Statistics are given in Table 10). For Cora/Pubmed, nodes are papers (title+abstract), edges are citations, and classes are research topics. For airports, nodes are airports, edges are flight connections, and classes correspond to activity levels (Ribeiro et al., 2017).
Baselines.
For text-attributed graphs, we compare Raw Feat. (bag-of-words/TF-IDF), Raw Text (use original text), TAPE (He et al., 2024), KEA(Chen et al., 2024b), and TANS (Wang et al., 2025a). For text-free graphs, we compare hand-crafted topology features—Node Degree, Eigenvector (Dwivedi et al., 2023), Random Walk (Dwivedi et al., 2022), and TANS Wang et al. (2025a). In all text-attributed baselines, generated texts are appended to the original node text and encoded by the same sentence encoder for fairness, while for text-free graphs, the generated texts are used directly as node descriptions.
Evaluation Protocol.
Unless otherwise specified, we focus on node classification with a GCN backbone (Kipf and Welling, 2017). We also report results with GAT (Veličković et al., 2018) and MLP in the text-attributed setting, following Wang et al. (2025a). For single-graph learning, we adopt the low-label / high-label splits: in the low-label regime, we use 20/30 nodes per class for train/valid on Cora/Pubmed (10/20 for Brazil); in the high-label regime, we use a 60/20/20 train/valid/test split. All reported numbers are averages over 30 random seeds with mean $\pm$ standard deviation, selecting models by the best validation accuracy. For the text encoder, we adopt MiniLM (Wang et al., 2020) for fair comparison with other methods, unless otherwise noted.
@@
Raw Text $76.97\pm 1.95$ $75.50\pm 2.03$ $70.78\pm 2.00$ $87.49\pm 0.54$ $87.20\pm 0.51$ $82.58\pm 0.38$ $4.33$
+ TAPE $76.50\pm 3.27$ $75.30\pm 1.92$ $71.06\pm 2.13$ $88.21\pm 0.62$ $87.80\pm 0.48$ $83.98\pm 0.59$ $3.83$
+ KEA $76.88\pm 1.73$ $75.74\pm 2.06$ $71.32\pm 2.51$ $88.10\pm 0.49$ $87.77\pm 0.50$ $85.33\pm 0.41$ $3.33$
+ TANS $76.27\pm 2.35$ $76.99\pm 2.02$ $73.64\pm 2.59$ $89.16\pm 0.39$ $7.98\pm 0.48 88$$8.84\pm 0.43$ $2.50$
+ GES (Ours) $79.52\pm 1.89$ $79.09\pm 1.50$ $76.51\pm 2.71$ $90.01\pm 0.52$ $88.42\pm 0.52$ $90.01\pm 0.52$ $1.00$
@@
Random Walk $56.70\pm 2.47$ $56.11\pm 2.11$ $69.70\pm 14.34$ $55.71\pm 4.01$ $62.80\pm 3.01$ $68.40\pm 9.65$ $4.00$
TANS $55.13\pm 1.52$ $60.61\pm 2.71$ $80.61\pm 12.14$ $56.33\pm 5.73$ $65.81\pm 3.11$ $71.60\pm 10.66$ $2.50$
GES (Ours) $\bm{56.80\pm 2.79}$ $\bm{61.66\pm 1.77}$ $80.91\pm 11.10$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{59.51\pm 4.44}$} $\bm{68.14\pm 2.41}$} $\bm{75.19\pm 7.42}$ $\bm{1.00}$
@@
3.2 Main Results
3.2.1 Results on Text-Attributed Graphs
Table 1 reports node classification accuracy on text-attributed Cora and Pubmed under both low- and high-label settings, using GCN, GAT, and MLP backbones. On these datasets, augmenting Raw Text with existing LLM-based methods consistently improves performance over Raw Feat. and Raw Text across all backbones in both label regimes. GES further improves upon the other methods in all configurations, consistently achieving higher accuracy than TANS for all backbones. These results indicate that exemplar-guided, structure-aware refinement is robust across architectures and scales of text-attributed citation graphs.
@@ 3.2.1 Results on Text-Attributed Graphs
3.2.2 Results on Text-Free Graphs
We evaluate GES on three text-free airport graphs using GCN as the backbone. Table 2 reports results under both low- and high-label splits. In the low-label regime, structural baselines such as Node Degree, Eigenvector, and Random Walk outperform the one-hot feature baseline, confirming the importance of topology. GES matches or surpasses these baselines on all three graphs, achieving the best accuracy across all datasets. Notably, although no human-written text is available, GES leverages structural cues to retrieve exemplars and induces task-aligned node semantics through exemplar-guided refinement. Overall, these results indicate that structure and semantic-aware refinement is particularly effective when raw node features are absent and structural roles dominate.
3.2.3 Results under Low Label Budget
We investigate how GES performs relative to TANS as the number of training labels per class varies. Because GES descriptions are generated through exemplar retrieval from the model-conditioned memory, they encode implicit class structure even before the GNN is trained on the target labels. When labeled data is abundant, the GNN has sufficient supervised signal to learn good representations regardless of description quality; when labels are scarce, the quality of input features becomes critical, and GES descriptions effectively provide soft supervision through the text itself. Figure 3 confirms this: at 10 labels/class, GES outperforms TANS on all five datasets (avg. +2.16%), with gains of +2.90% on Pubmed and +4.06% on Brazil. At extreme scarcity (5 labels/class), GES wins on 4 of 5 datasets, with the sole exception being Cora, where rich text semantics already provide strong discriminative signal. These results indicate that GES is most valuable in the realistic low-resource regime where labeled graph data is expensive to obtain.
Figure 3: Label budget sensitivity across five datasets.
3.2.4 Results under Domain Adaptation
We assess cross-graph generalization in domain adaptation, where the model is trained on a source graph and directly evaluated on a target graph without fine-tuning. We train on a source graph and evaluate on a target graph (20% val / 80% test on target). We adopt the source-target pairs from Wang et al. (2025a) on the three airport graphs. Results in Table 3 show that TANS already improves over SVD-based alignment and purely structural features in most transfer directions. GES further increases accuracy in the majority of source-target pairs. For example, in USA to Europe or Brazil, and from Europe to Brazil, we achieve the best performance among all methods. The performance gains of GES are especially notable on more challenging transfers involving Brazil, where structural roles differ markedly across graphs. These improvements suggest that structure and semantic-aware node texts provide a more transferable representation than synthesized descriptions.
3.2.5 Results under Transfer Learning
@@
Raw Feat. + SVD $70.39\pm 6.12$ $70.48\pm 3.71$
Raw Text $75.77\pm 2.96$ $79.62\pm 2.04$
+ TAPE $75.60\pm 2.39$ $79.25\pm 2.06$
+ KEA $75.25\pm 2.50$ $79.59\pm 1.61$
+ TANS $76.14\pm 2.28$ $80.05\pm 1.74$
  + GES $79.80\pm 1.76$ $80.33\pm 1.57$
Table 4: Pretrain–finetune transfer (GCN, low-label). C = Cora, P = Pubmed.
[…]
4 Analysis
This section analyzes the key mechanisms underlying GES beyond aggregate performance. We focus on how exemplar retrieval and semantic refinement shape node representations. We provide additional analysis in Appendix F and G.
@@ 4 Analysis
Retrieval Cora USA
GES (Ours) 89.31 $\pm$ 1.14 68.14 $\pm$ 2.41
@@
Random 89.06 $\pm$ 0.92 66.47 $\pm$ 3.24
Table 5: Ablation on retrieval strategy.
@@
We examine the role of the model-conditioned memory in exemplar retrieval for semantic refinement. To isolate its effect, we compare GES with three ablated variants: Random exemplar selection, Structure-only retrieval based solely on structural similarity, and Text-only retrieval based solely on semantic similarity. All variants use the same refinement procedure and prompt format. Results in Table 5 show that joint semantic–structural retrieval in GES consistently outperforms all ablated variants on both Cora and USA. On Cora, text-only retrieval performs competitively, reflecting the strong semantic signal in raw node texts, while structure-only retrieval is weaker. In contrast, on the text-free USA graph, both single-modality variants degrade performance, indicating that neither modality alone is sufficient for stable refinement.
Sensitivity to Retrieval Hyperparameters.
Figure 4: Sensitivity to $K$, $\tau$, and $\alpha$. Stars mark the best setting per dataset.
GES introduces three retrieval hyperparameters: the support-set size $K$, the entropy threshold $\tau$, and the semantic–structural trade-off $\alpha$ in $S(v,u)=\alpha\,\mathrm{sim}_{t}(v,u)+(1{-}\alpha)\,\mathrm{sim}_{s}(v,u)$. Figure 4 reports accuracy on Cora and USA (high-label, GCN) as each parameter is varied while the others are held at their defaults. Performance is stable across a wide range of all three parameters. Higher $\alpha$ and $\tau$ is preferred on text-rich Cora, while lower $\alpha$ and $\tau$ favors the text-free USA, consistent with the underlying structure–semantics regime.
Robustness to LLM Choice.
Dataset TANS GES
Cora $81.79\pm 1.21$ $\mathbf{82.00\pm 1.51}$
Pubmed $83.00\pm 1.98$ $\mathbf{83.80\pm 1.79}$
USA $60.61\pm 3.23$ $\mathbf{61.69\pm 2.41}$
Europe $56.40\pm 3.64$ $\mathbf{56.45\pm 3.21}$
Brazil $73.94\pm 10.43$ $\mathbf{75.45\pm 13.92}$
Table 6: GES vs. TANS with Gemma-3-12B.
Table 6 shows that replacing GPT-4o-mini with the open-weight Gemma-3-12B (Team, 2025) in the low-label setting preserves the improvement of GES over TANS on all four datasets. The gains are largest on structure-dominated graphs, where exemplar-guided refinement adds the most value regardless of LLM capacity. This confirms that the improvements stem from the retrieval and refinement mechanism itself, not from the specific language model, and that GES can be deployed with open-weight models in resource-constrained settings.
Cost–Accuracy Tradeoff.
[…]
Cora +2.3M +0.14
Pubmed +16.8M +0.84
USA +1.0M +2.33
Europe +0.3M +3.18
Brazil +0.1M +3.59
Table 7: Token overhead vs. accuracy gain of GES over TANS.
A key practical question is whether the gains of GES justify its additional inference cost. GES issues exactly one LLM call per node; the only overhead over other methods is the inclusion of retrieved exemplars in the prompt (${\sim}850$ additional input tokens per node). Table 7 reports the token overhead of GES over TANS. On Pubmed, the largest graph (19.7k nodes), this amounts to 16.8M extra tokens—substantial in absolute terms, but yielding +0.84% in the high-label setting and +3.25% in the low-label setting (Table 1 and Table 7). On airport graphs, all three datasets combined require only 1.4M extra tokens, yet GES achieves +2.33% on USA, +3.18% on Europe, and +3.59% on Brazil. GES requires a single GNN pre-training pass to populate the memory, which is negligible relative to LLM inference.
@@ 5 Related Work
GNNs and Fixed Inductive Bias. Classical GNNs learn representations through local message passing, thereby encoding fixed inductive biases such as locality and homophily (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Xu et al., 2019). Structure-only methods based on random walks or structural roles further impose predefined topological assumptions (Perozzi et al., 2014; Grover and Leskovec, 2016; Ribeiro et al., 2017). While effective in specific regimes, these methods rely on fixed priors that do not adapt to heterogeneous structure–semantics distributions across graph domains.
@@
Model-Centric Refinement. Most existing attempts to address structure–semantics heterogeneity remain model-centric. This includes strengthening GNN architectures with richer structural priors (Maron et al., 2018; Murphy et al., 2019; Jin et al., 2022), treating LLMs as direct graph reasoners via graph-to-text serialization (Zhao et al., 2023a; Kong et al., 2024; Chen et al., 2024a), and introducing auxiliary models to enrich node representations with external semantic signals (Yao et al., 2019; He et al., 2023; Yang et al., 2021). Despite their diversity, these approaches all inject additional inductive biases from the model side, and the construction of node representations remains largely static once the model is specified.
@@
Data-Centric Refinement and Our Perspective. Existing data-centric methods primarily focus on graph augmentation, structure learning, or pseudo-labeling for robustness and generalization (You et al., 2020; Jin et al., 2020; Chen et al., 2023; He et al., 2023; Chen et al., 2024b), rather than on adapting node semantics to task-specific structure–semantics regimes. In contrast, our work treats node semantics as task-adaptive variables and proposes a framework for exemplar-guided, structure-aware semantic refinement, providing a fundamentally different data-centric perspective for handling structure–semantics heterogeneity.
The Position of Our Work.
In contrast to the above data-centric paradigms, our work targets a fundamentally different objective. Rather than augmenting data for invariance, modifying graph structure, or propagating pseudo-label supervision, we focus on task-driven refinement of node semantics themselves. Our method treats node semantics as adaptive variables that are reshaped under structural context and predictive feedback from a downstream GNN, enabling direct handling of structure–semantics heterogeneity at the level where the balance between structure and semantics is instantiated.
6 Conclusion and Limitations